# DART: Distilling Autoregressive Reasoning to Silent Thought

### v1

@@ Abstract
Chain-of-Thought (CoT) reasoning has significantly advanced Large Language Models (LLMs) in solving complex tasks. However, its autoregressive paradigm leads to significant computational overhead, hindering its deployment in latency-sensitive applications. To address this, we propose DART (Distilling Autoregressive Reasoning to Silent Thought), a self-distillation framework that enables LLMs to replace autoregressive CoT with non-autoregressive Silent Thought (ST). Specifically, DART introduces two training pathways: the CoT pathway for traditional reasoning and the ST pathway for generating answers directly from a few ST tokens. The ST pathway utilizes a lightweight Reasoning Evolvement Module (REM) to align its hidden states with the CoT pathway, enabling the ST tokens to evolve into informative embeddings. During inference, only the ST pathway is activated, leveraging evolving ST tokens to deliver the answer directly. Extensive experimental results demonstrate that DART achieves comparable reasoning performance to existing baselines while offering significant efficiency gains, serving as a feasible alternative for efficient reasoning.
@@ 3.1 Dual-Pathway Architecture
Notably, the first $t-1$ tokens of $Z$ are indeed CoT tokens, while the remaining are separator tokens shared with the ST pathway. In this paper, we fix $z_{t:N}$ as the answer prompt "Answer:".
@@ 4 Experiments
To validate the design of DART, we conduct extensive experiments and present the key results. Additional implementation details and extended findings are provided in Appendices B-D.
@@ 4.1 Experimental Settings
Datasets. Following previous work Deng et al. (2024); Hao et al. (2024), we fine-tune models on GSM8K-Aug (Deng et al., 2023), an augmented version of GSM8K (Cobbe et al., 2021), which includes diverse reasoning traces. To mitigate the risk of models memorizing the final answer from CoT sequences, the last step is omitted during training Shen et al. (2025). For out-of-distribution evaluation, we adopt GSM-HARD (Gao et al., 2023), SVAMP (Patel et al., 2021), and MultiArith (Roy and Roth, 2015) as robustness benchmarks.
@@ 4.2 Empirical Results
Comparison between Baselines. Table 1 summarizes the performance across GSM8K, GSM-HARD, SVAMP, and MultiArith. To ensure a fair comparison of efficiency, we measure the inference time of all baselines on a Nvidia A10 GPU, even for those whose accuracy is sourced from prior reports. As shown, DART achieves the best performance among all NAR baselines on GSM8K, delivering a notable 10.1% accuracy gain with negligible latency overhead (only 1 ms). On all out-of-distribution datasets, DART consistently outperforms other NAR methods, indicating robust generalization beyond the training distribution. While AR methods obtain higher accuracy, they suffer from significantly reduced inference efficiency due to stepwise generation, despite efforts to compress reasoning steps. These results demonstrate that DART achieves a compelling trade-off between accuracy and efficiency by fully leveraging distilled CoT knowledge in a single-step latent space.
@@
Ablation Study. To validate the contributions of key DART components, we evaluate the variants with certain components omitted or replaced. Our findings in Table 2 are as follows: (1) Omitting $\mathcal{L}_{distill}$, which transfers reasoning patterns from CoT trajectories, leads to a substantial performance drop; (2) Excluding the ST tokens impairs accuracy, likely due to the loss of CoT-derived positional priors; (3) Our proposed REM significantly enhances reasoning capability compared to using either no additional module or the vanilla LoRA.
4.3 Interpretability Analysis
@@
We decode the final hidden states of ST tokens into natural language using the model’s word embeddings, finding that 69.9% of the translated tokens match the words in ground-truth CoT. This ratio further rises to 80.1% for cases that yield the correct answer. Additionally, we observe that the ST tokens also reflect a progressive reasoning process similar to CoT. For example, the top-1 ST translation is "192 192 192 192 24 24 24 24 24 24 24 24 24 24 24 24 24 24 24 24", corresponding to the CoT "<<3*64=192>> <<192/8=24>>".
@@ 5 Conclusion
We present DART, a fine-tuning framework that empowers LLMs to perform implicit reasoning in a non-autoregressive manner. By distilling knowledge from CoT data, the models trained with DART achieve a remarkable balance between accuracy and latency using the evolving ST tokens. Extensive experiments confirm DART’s robustness and the effectiveness of its design.
@@ Appendix A Analysis of the Shift Value
[…]
@@
where $W_{K}^{l},W_{V}^{l}\in\mathbb{R}^{n\times n}$ are the key and value projection matrices of the $l$-th decoder layer, $H_{Q}$, $H_{Z}$ are the input hidden state of question $Q$ and intermediate tokens $Z$, $\mathbf{q}^{l}$ is the attention query vector corresponding to $z_{N}$, and $\mathbf{a}^{l}$ is the activation when only $Q$ is given. The superscript $l-1$ for $H_{Q}$, $H_{Z}$ is omitted for simplicity. The approximation in the second step is obtained by omitting the softmax operation and scaling factor $\sqrt{n}$. Then, by going through the feed-forward layer $f(\cdot)$, we can get the hidden state of $z_{N}$ in the $l$-th layer as follows:
@@ Appendix C Implementation Details
For all experiments, we set $d=128$ and $\alpha=32$ for REM, consistent with the configuration used for LoRA (Hu et al., 2022) to fine-tune the CoT pathway. We employ the AdamW (Loshchilov and Hutter, 2019) with a cosine annealing learning rate schedule following a 3% warm-up period. Additionally, we enable the bf16 in the trainer and evaluate the models using bfloat16 precision.
We fine-tune the Llama-3.2-1B, the primary base model in our experiments, for 10 epochs with the learning rate initialized as 8e-4.
@@
Appendix D More Experimental Results.
Sensitivity Analysis on ST Token Number. We conduct the experiments on GSM8K with various values of the ST token number $C$. As shown in Figure 2, the accuracy rapidly increases during the initial stage as $C$ grows, then stabilizes once the token number becomes sufficient, demonstrating both the necessity and robustness of using ST tokens. Furthermore, thanks to the non-autoregressive paradigm, DART introduces negligible latency even as $C$ increases.
Experiments on Various Base Models. To further demonstrate the robustness of DART across different LLMs, we conduct the experiments using Qwen2.5-1.5B (Team, 2024) and GPT2 (Radford et al., 2019) as base models. For Qwen2.5-1.5B, we initialize the learning rate at 5e-4 and keep all other configurations as those used for Llama-3.2-1B. For GPT2, we set the initial learning rate to 2e-3 and train the model for 40 epochs. As shown in Table 5, DART consistently boosts the reasoning capabilities without notable latency.
GPT2 Qwen2.5-1.5B
Methods Acc IT Acc IT
No-CoT 16.2 18 33.3 95
DART 24.7 19 42.2 95
Table 5: Results on GPT2 and Qwen2.5-1.5B. Acc and IT indicate the Accuracy (%) and inference time (ms, on a Nvidia A10 GPU), respectively.
Methods Accuracy
No-CoT 32.5
DART (w/ $\mathcal{L}_{distill}$ on $z_{N}$) 42.6
      w/ $\mathcal{L}_{distill}$ on $y_{1}$ 31.5
      w/ $\mathcal{L}_{distill}$ on $[z_{N};Y]$ 33.7
Table 6: Accuracy (%) with various tokens used in $\mathcal{L}_{distill}$.
Different Choices of Alignment Token. To empirically demonstrate the rationality of applying $\mathcal{L}_{distill}$ on $z_{N}$, we perform the experiments with different alignment tokens. As shown in Table 6, employing other tokens like the answer tokens in $\mathcal{L}_{distll}$ hinders the effectiveness of alignment.

### v2

@@ Abstract
Chain-of-Thought (CoT) reasoning has significantly advanced Large Language Models (LLMs) in solving complex tasks. However, its autoregressive paradigm leads to significant computational overhead, hindering its deployment in latency-sensitive applications. To address this, we propose DART (Distilling Autoregressive Reasoning to Silent Thought), a self-distillation framework that enables LLMs to replace autoregressive CoT with non-autoregressive Silent Thought (ST). Specifically, DART introduces two training pathways: the CoT pathway for traditional reasoning and the ST pathway for generating answers directly from a few ST tokens. The ST pathway utilizes a lightweight Reasoning Evolvement Module (REM) to align its hidden states with the CoT pathway, enabling the ST tokens to evolve into informative embeddings. During inference, only the ST pathway is activated, leveraging evolving ST tokens to deliver the answer directly. Extensive experimental results demonstrate that DART offers significant performance gains compared with existing non-autoregressive baselines without extra inference latency, serving as a feasible alternative for efficient reasoning.
@@ 3.1 Dual-Pathway Architecture
Notably, the first t1t-1 tokens of $Z$ are indeed CoT tokens, while the remaining are separator tokens shared with the ST pathway. In this paper, we fix $z_{t:N}$ as the answer prompt "Answer:".
@@ 3.2 REM-based Self-Distillation
Methods FLOPs peak GPU memory
CoT 1874774802 12.28
No-CoT 1398731734 11.70
DART 1881308345 18.07
Table 2: Total FLOPs (GF) and peak GPU memory (GB) for CoT, No-CoT, and DART. The reported peak GPU memory is the average across 8 Nvidia A10 GPU
@@ 4 Experiments
To validate the design of DART, we conduct extensive experiments and present the key results. Additional implementation details are provided in Appendices B-C.
@@
      w/o $\mathcal{L}_{distill}$ 33.7
      w/ $\mathcal{L}_{distill}$ on $y_{1}$ 31.5
      w/ $\mathcal{L}_{distill}$ on $z_{N};Y]$ 33.7
@@ 4.1 Experimental Settings
Datasets. Following previous work Deng et al. (2024); Hao et al. (2024), we mainly fine-tune models on the complicated mathematical reasoning benchmark GSM8K-Aug (Deng et al., 2023), an augmented version of GSM8K (Cobbe et al., 2021), which includes diverse math reasoning traces. To mitigate the risk of models memorizing the final answer from CoT sequences, the last step is omitted during training Shen et al. (2025). For out-of-distribution evaluation, we adopt GSM-HARD (Gao et al., 2023), SVAMP (Patel et al., 2021), and MultiArith (Roy and Roth, 2015) as robustness benchmarks.
@@ 4.2 Main Results
Comparison between Baselines. Table 1 summarizes the performance across GSM8K, GSM-HARD, SVAMP, and MultiArith. To ensure a fair comparison of efficiency, we measure the inference time of all baselines on a Nvidia A10 GPU, even for those whose accuracy is sourced from prior reports. As shown, DART achieves the best performance among all NAR baselines on GSM8K, delivering a notable 10.1% accuracy gain with negligible latency overhead (only 1 ms). On all out-of-distribution datasets, DART consistently outperforms other NAR methods, indicating robust generalization beyond the training distribution. While AR methods obtain higher accuracy, they suffer from significantly reduced inference efficiency due to stepwise generation, despite efforts to compress reasoning steps. These results demonstrate that DART achieves a compelling trade-off between accuracy and efficiency by fully leveraging distilled CoT knowledge in a single-step latent space. For training-compute transparency, we provide the total FLOPs and peak GPU memory for CoT, No-CoT, and DART in Table 2. All experiments were conducted on 8 Nvidia A10 GPUs. We can observe that DARt shares a comparable total FLOPs with the CoT baseline. Moreover, it’s important to note the efficiency of DART’s trainable parameters. For Llama-3.2-1B, CODI uses 98,574,336 trainable parameters, and Coconut employs a more complex multi-stage fully fine-tuning. In contrast, DART only has 44,040,192 trainable parameters, demonstrating an excellent efficiency trade-off considering both training and inference costs.
@@
Ablation Study. To validate the contributions of key DART components, we evaluate the variants with certain components omitted or replaced. Our findings in Table 3 are as follows: (1) Omitting $\mathcal{L}_{distill}$, which transfers reasoning patterns from CoT trajectories, leads to a substantial performance drop; (2) Excluding the ST tokens impairs accuracy, likely due to the loss of CoT-derived positional priors; (3) Our proposed REM significantly enhances reasoning capability compared to using either no additional module or the vanilla LoRA; (4) Employing other tokens like the answer tokens as the distilling token in $\mathcal{L}_{distill}$ hinders the effectiveness of alignment, empirically demonstrating the rationality of applying $\mathcal{L}_{distill}$ on $z_{N}$.
Sensitivity Analysis on ST Token Number. We conduct the experiments on GSM8K with various values of the ST token number $C$. As shown in Figure 2, the accuracy rapidly increases during the initial stage as $C$ grows, then stabilizes once the token number becomes sufficient, demonstrating both the necessity and robustness of using ST tokens. Furthermore, benefiting from the non-autoregressive paradigm, DART introduces negligible latency even as $C$ increases.
4.3 Further Analysis
ProsQA CommonsenseQA
Methods Acc IT Acc IT
CoT 98.8 882 68.5 1471
No-CoT 94.0 129 65.4 34
DART 99.4 130 75.1 34
Table 4: Results on ProsQA and CommonsenseQA-CoT. Acc and IT indicate the Accuracy (%) and inference time (ms, on a Nvidia A10 GPU), respectively.
Task Generalizability. To evaluate generalizability in other reasoning tasks, we fine-tune the models on the logic reasoning benchmark ProsQA (Hao et al., 2024) and commonsense reasoning benchmark CommonsenseQA-CoT Shen et al. (2025). As summarized in Table 4, DART can even outperform the CoT baseline in ProsQA and CommonsenseQA-CoT while maintaining its low inference latency, demonstrating excellent robustness across diverse reasoning domains.
Robustness across Various Models. To further demonstrate the robustness of DART across different LLMs, we conduct the experiments using Llama-3.2-3B, Qwen2.5-1.5B (Team, 2024) and GPT2 (Radford et al., 2019) as base models. Experimental results in Table 5 show that DART can consistently boost the reasoning capabilities without notable latency across these models, demonstrating its scalability to larger and smaller models.
No-CoT DART
Models Acc IT Acc IT
GPT2 16.2 18 24.7 19
Qwen2.5-1.5B 33.3 95 42.2 95
Llama-3.2-3B 40.8 62 46.6 62
Table 5: Results on GPT2, Qwen2.5-1.5B and Llama-3.2-3B. Acc and IT indicate the Accuracy (%) and inference time (ms, on a Nvidia A10 GPU), respectively.
@@ 4.3 Further Analysis
Qualitative Analysis. We decode the final hidden states of ST tokens into natural language using the model’s word embeddings, finding that 69.9% of the translated tokens match the words in ground-truth CoT. This ratio further rises to 80.1% for cases that yield the correct answer. Additionally, we observe that the ST tokens also reflect a progressive reasoning process similar to CoT. For example, the top-1 ST translation is "192 192 192 192 24 24 24 24 24 24 24 24 24 24 24 24 24 24 24 24", corresponding to the CoT "<<3*64=192>> <<192/8=24>>".
@@ 5 Conclusion
We present DART, a fine-tuning framework that empowers LLMs to perform implicit reasoning in a non-autoregressive manner. By distilling knowledge from CoT data, the models trained with DART achieve a remarkable balance between accuracy and latency using the evolving ST tokens. Furthermore, extensive experiments confirm DART’s robustness across various benchmarks and validate the effectiveness of its design.
@@ Appendix A Analysis of the Shift Value
$\left[\cdot;\cdot\right]$ The concatenation operation for matrix pairs and sequence pairs.
@@
where $W_{K}^{l},W_{V}^{l}\in\mathbb{R}^{n\times n}$ are the key and value projection matrices of the $l$-th decoder layer, $H_{Q}$, $H_{Z}$ are the input hidden state of question $Q$ and intermediate tokens $Z$, $\mathbf{q}^{l}$ is the attention query vector corresponding to $z_{N}$, and $\mathbf{a}^{l}$ is the activation when only $Q$ is given. The superscript l1l-1 for $H_{Q}$, $H_{Z}$ is omitted for simplicity. The approximation in the second step is obtained by omitting the softmax operation and scaling factor $\sqrt{n}$. Then, by going through the feed-forward layer $f(\cdot)$, we can get the hidden state of $z_{N}$ in the $l$-th layer as follows:
@@ Appendix B Datasets
ProsQA Question = "Every yimpus is a yumpus. Jack is a yerpus. Max is a vumpus. Max is a zhorpus. Every brimpus is a rorpus. Every brimpus is a timpus. Every rorpus is a hilpus. Every gwompus is a yumpus. Every gorpus is a lempus. Every impus is a kerpus. Every impus is a brimpus. Every impus is a hilpus. Every kerpus is a boompus. Max is a gorpus. Every gorpus is a rempus. Every gorpus is a yimpus. Every gorpus is a rompus. Every yerpus is a timpus. Every kerpus is a rorpus. Every yerpus is a impus. Every gwompus is a yimpus. Every gorpus is a yumpus. Max is a gwompus. Every brimpus is a kerpus. Every gwompus is a vumpus. Every yumpus is a gerpus. Is Jack a zhorpus or boompus?" CoT = "Jack is a yerpus. Every yerpus is a impus. Every impus is a kerpus. Every kerpus is a boompus." Answer = "Jack is a boompus."
CommonsenseQA-CoT Question = "The fox walked from the city into the forest, what was it looking for? Choices: A: pretty flowers. B: hen house C: natural habitat D: storybook E: dense forest" CoT = "The fox, being a wild animal, would typically seek its natural habitat when moving from the city into the forest. Options like "pretty flowers," "hen house," and "storybook" do not align with a fox’s natural behavior, while "dense forest" describes the environment but not the fox’s purpose. Therefore, "natural habitat" is the most logical choice." Answer = "C"
@@ Appendix C Implementation Details
For all experiments, we set $d=128$ and $\alpha=32$ for REM, consistent with the configuration used for LoRA (Hu et al., 2022) to fine-tune the CoT pathway. We find that the distillation loss is much smaller than the other two losses. We set the trade-off hyperparameter $\lambda=20$ since we find that the distillation loss is much smaller than the other two losses. We employ the AdamW (Loshchilov and Hutter, 2019) with a cosine annealing learning rate schedule following a 3% warm-up period. Additionally, we enable the bf16 in the trainer and evaluate the models using bfloat16 precision.
We fine-tune the Llama-3.2-1B, the primary base model in our experiments, for 10 epochs with the learning rate initialized as 8e-4. For Llama-3.2-3B and Qwen2.5-1.5B, we initialize the learning rate at 5e-4 and keep all other configurations as those used for Llama-3.2-1B. For GPT2, we set the initial learning rate to 2e-3 and train the model for 40 epochs.